# Lexical categories of stem-forming roots in Mapudüngun verb forms

### v2

@@ Abstract
After developing a computational system for morphological analysis of the Mapuche language, and evaluating it with texts from various authors and styles, it became necessary to verify the linguistic assumptions of the source used as the basis for implementing this tool.
In the present work, the primary focus is on the lexical category classification of Mapudüngun roots recognised as verbal in the source utilised for the development of the morphological analysis system.
The results of this lexical category revision directly benefit the computational analyser, as they are implemented as soon as they are verified. Additionally, it is hoped that these results will help clarify some uncertainties about lexical categories in the Mapuche language.
This work addresses a preliminary task to identify the valency of true verbal roots, the results of which will be presented in a subsequent work that complements this article.
Resumen
@@
Tras haber desarrollado un sistema computacional para el análisis morfológico de la lengua mapuche y de evaluarlo con textos de diversos autores y estilos, se ha visto la necesidad de verificar los postulados lingüísticos de la fuente utilizada como base para implementar esta herramienta.
En el presente trabajo, el enfoque principal es la clasificación de la categoría léxica de las raíces del mapudüngun reconocidas como verbales en la fuente utilizada para el desarrollo del sistema de análisis morfológico.
Los resultados de esta revisión categorial benefician directamente al analizador computacional, ya que se implementan tan pronto como se verifican. Además, se espera que estos resultados ayuden a aclarar algunas incertidumbres sobre las categorías léxicas en la lengua mapuche.
Este trabajo aborda una tarea preliminar a la identificación de la valencia de las raíces verbales, cuyos resultados se presentarán en un trabajo posterior que complementa este artículo.
Resum
Després d’haver desenvolupat un sistema computacional d’anàlisi morfològica de la llengua mapuche, i d’avaluar-lo amb textos de diversos autors i estils, es va veure la necessitat de verificar els postulats lingüístics de la font utilitzada com a base per implementar aquesta eina.
En el present treball, l’enfocament principal és la classificació de la categoria lèxica de les arrels del mapudüngun reconegudes com a verbals en la font utilitzada per al desenvolupament del sistema d’anàlisi morfològica.
Els resultats d’aquesta revisió categorial beneficien directament l’analitzador computacional, ja que s’implementen tan bon punt es verifiquen. A més, s’espera que aquests resultats ajudin a aclarir algunes incerteses sobre les categories lèxiques en la llengua mapuche.
Aquest treball aborda una tasca preliminar a la identificació de la valència de les arrels verbals, els resultats de la qual es presentaran en un treball posterior que complementa aquest article.
@@
Keywords: Mapuche; Mapudüngun; FST; finite state transducer; morphology; morphosyntax; affixes; suffixes; morphemes; verbs; valency; verbal roots; parts of speech;
@@
Palabras clave: Mapuche; Mapudüngun; FST; transductor de estados finitos; morfología; morfosintaxis; afijos; sufijos; morfemas; verbos; valencia; raíces verbales; categorías gramaticales; categorías léxicas;
@@
Paraules clau: Maputxe; Mapudüngun; FST; transductor d’estats finits; morfologia; morfosintaxi; afixos; sufixos; morfemes; verbs; valència; arrels verbals; categories gramaticals; categories lèxiques;
@@ 1 Introduction
This article details the methods used for the reclassification of the lexical roots involved in Mapuche verb themes and presents the results obtained from this process. This task is essential to accurately determine the valency (in a subsequent work, Chandía, 2025) of roots that are genuinely verbal. Without this step, there is a risk of incorrectly attributing verbal properties to roots that do not belong to this lexical category.
@@
The task is described as a reclassification because it involves revisiting the lexical category classification of roots found in Mapudüngun verb themes (see §2.1, p. 2.1) which was originally established by the anthropologist Ineke Smeets [^1] This author will be cited continuously for her 2008 work; thereafter, only her last name and the referenced page will be provided, omitting the year.
in her work on the grammatical description of the Mapuche language, published in 2008 with the title ‘‘A Grammar of Mapuche’’.
@@
This research, like others conducted within the Düngupeyüm [^2] Düngupeyüm is the name of the computational morphological analysis system that has been implemented; the name derives from düngu- ‘language, speech, concern’, -pe- suffix indicating physical or temporal proximity, -yü- suffix indicating a habitual aspect feature, and the affix -m, a nominaliser indicating place or instrument. The translation would be something like ‘instrument always used for language’, i.e., ‘language tool’. The system is publicly available on https://www.chandia.net/dungupeyum
project, relies on the grammatical description provided by Smeets. Our goal is to enhance the computational morphological analysis system that we have developed using finite-state transducers (Chandía, 2012, 2021, 2022), which initially implemented the anthropologist’s work. This focus on the elements and dynamics of the language allows us to refine and expand the capabilities of the system.
@@
For a computational system such as the one described earlier, accurately determining both the lexical category of roots and the valency of verbal roots is crucial for proper analysis. This is because, in Mapudüngun, the addition of suffixes to verbal forms, particularly aspectual ones, depends on the verbal valency or its modification during word formation. Consequently, verbal valency plays a key role in analysing the form to identify the implicated suffixes. Therefore, it was necessary to verify the lexical category of roots involved in the formation of verbal themes, as Mapudüngun naturally verbalises roots from other lexical categories, including nouns, adjectives, adverbs, and even some pronouns and numerals. The aim is to accurately recognise the valency of roots that are genuinely verbal, avoiding misclassification of roots belonging to other categories. This document focuses on the task of lexical classification, while the classification of valency will be addressed in a subsequent study (Chandía, 2025).
@@
During the extensive period dedicated to developing the morphological analysis tool, several observations made by Smeets needed to be verified, and some required modification. She provided a grammatical description of Mapudüngun based on a group of four informants [^3] As asserted by Smeets in email correspondence, the number may be increased by four or five, which where sporadic informants, thus bringing the total to approximately ten.
. As a result, her work does not fully capture the entirety of the Mapuche language or all the phenomena present within it.
@@
The main contribution of Smeets’ ‘‘A Grammar of Mapuche’’, which is particularly valuable for the development of a computational morphological analysis system, is the identification of a prevalent or fixed position for approximately one hundred suffixes that can appear in Mapuche verbal forms. Additionally, she provides the morphotactics [^4] Morphotactics refers to the rules governing the allowed and prohibited combinations of morphemes (both among themselves and with roots) in a language. These rules influence certain forms within the language. For example, in a Mapuche nominal form, morphemes indicating verbal mode cannot appear unless the nominal form has been verbalised. Similarly, in a verbal form, a nominaliser affix and a grammatical person morpheme cannot co-occur. Morphotactic rules can be influenced by grammatical, semantic and morphological factors.
, or combinatorial rules, for most of these suffixes.
@@
Smeets states that suffixes appear in relatively fixed positions along the verbal form. Based on their placement relative to other suffixes and their function, these verbal affixes are organised into thirty-six numbered slots, arranged from right to left. Slot one is at the far right end of the form, while slot thirty-six is next to the root on the left. Certain slots contain suffixes that are mutually exclusive for grammatical or semantic reasons.
@@
E.01)
Verb with 10 suffixes (Smeets: 443 .76)
nünieñmarputueyiñmu
  nü            -nie      -ñma  -r      -pu     -tu    -e      -y     -ø   -iñ    -mu
-TV [^5] The meaning of the labels and abbreviations can be found in Appendix §4.1.
.tomar +PRPS +IO   +ITR +LOC +RE +INV +IND +1 +PL +3A
‘they continued to take it away from us’ – ‘continuaron quitándonosla’,
lit.: ‘there they went continuously taking it away from us’ [^6] The Spanish translations of the examples from Smeets, which she translated into English, have been added. Literal interpretations of the Mapuche phrases have also been made in most cases, in an attempt to clarify the meaning of the phrase in the original language.
Figure 1: Graphical representation of the slots in Mapuche verbal forms (root and suffixes) Minimal intransitive verb form of 2nd person plural S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . X X X . . . Minimal transitive verb form of 2nd → 1st persons, plural S . . . . . . . . . . . X . . . . . . . . . . . . . . . . . . X ø X . . . Representation of example E.01 S . . . . X . . . . . X . . . . . . . X X X . . . . . . . . . X . X ø X X
@@
In the Mapuche language, a verbal theme [^7] To simplify, the term “verbal theme” refers to any root, whether simple or complex, to which suffixes are added to create a complete verbal predication. A simple “verbal theme” consists of a single root, while a complex “verbal theme” may involve two roots, a root with some suffixes, or a combination of these elements. Although it could be argued that these are more accurately described as lemmas, they are all referred to as “verbal theme” for the sake of simplicity (Chandía, 2021: 3).
can consist of one, two, or occasionally three roots from different lexical categories, which may combine with each other and sometimes with certain suffixes. A verbal theme can be simple, containing just one root from any lexical category (examples E.02.a, E.02.b, E.02.c), or complex, with several possible configurations:
    A root combined with one or more suffixes (example E.02.d), sometimes up to three suffixes.
@@
      pichi          -ka        -el (Smeets: 76 .16)
    -AJ.little +CONT +OVN
@@
      tripa         -y      -ø (Smeets: 67 .51)
    -IV.to.leave +IND +3
@@
    -IV.to.come +CA +IND1SG
    ‘I brought’; lit.:‘I made come’
@@ 1.1.1 Dictionaries
To date, the most reliable collections of words classified by their lexical type are the dictionaries published throughout the relatively short written history of the Mapudüngun language, particularly the following three [^8] All three dictionaries are accessible on the CORLEXIM site: https://www.chandia.net/corlexim
(Chandía, 2014).
@@
    Havestadt’s dictionary is a fourth lexical source occasionally consulted in cases of doubt. Published in 1777, it provides definitions in Latin [^9] Resource available for consultation via the following link: https://benmolineaux.github.io/bookshelf/1777HAVE-Vocab.html. An alternative Spanish version is also available at the following URL: https://benmolineaux.github.io/bookshelf/1777Havestadt-ParsQuarta-MAPU-SPA.html (Martinez and Carmona, 2023).
    and can be a valuable resource, particularly when dealing with variations in forms.
@@ 1.1.2 Computational Approach
AVENUE was the name of this initiative, which produced a document entitled ‘‘Informe Final’’ (Álvarez et al., 2005) summarising the project’s achievements and the works still to be done. When it started in 1999, it was known as NICE (Native-Language Interpretation and Communication Environment), later Avenue. Its goals were to create a simultaneous translation system and to contribute to the preservation and development of Indo-American languages [^10] Avenue worked with three indigenous languages, setting up teams in three countries: Aymara in Bolivia, Quechua in Peru and Mapudüngun in Chile.
. They were committed to creating their own alphabet for the language, transcribing texts according to this alphabet, extracting grammatical rules from the texts, compiling a parallel corpus of Mapudüngun-Spanish, also a glossary, developing a spell-checker of Mapudüngun for OpenOffice, and building a prototype of an automatic translator of Mapudüngun-Spanish. The report mentioned above states that these tasks were carried out at different levels of completion. The section dealing with the automatic translator mentions that it is an example-based machine translation system [^11] Example-based machine translation, or EBMT, works by searching its database for full translations or, more commonly, parts of sentences that have previously been translated by human translators.
. At the time of writing, however, researchers say they are working on a transfer-based machine translation system [^12] In a Transfer-Based Machine Translation (TBMT) system, the original text is first analysed morphologically and syntactically, resulting in a shallow syntactic representation. This representation is then transformed into a more abstract one that emphasises aspects relevant to the translation process and ignores other information. The transfer process converts this latter representation, still linked to the source language, into a representation at the same level of abstraction but linked to the target language. These two representations are called normalised or intermediate representations. From here the process is reversed: the syntactic components generate a representation of the text and finally the translation is generated.
in which the basic constructions of Mapudüngun have been recorded: simple sentences with intransitive and transitive verbs, noun phrases with determiners and modifiers, verb phrases with different temporal and aspectual specifications, passive verb phrases, and verb phrases with inverse agreement [^13] The inversion or reverse form in Mapudüngun is one of the possible expressions of the transitive verb (the others are direct and passive constructions) in which there is a third person acting upon a first or second person, this is expressed by the suffix -e before the mood suffix, and the suffix -(m)ew at the end, provided that the first or second person is a patient; i.e. the order of the constituents, which in Mapudüngun are suffixes, is not changed, but the detailed suffixes are added and combined to change the direction of the action of the verb (Zúñiga, 2006: 114)
. Also relevant is the section dealing with the morphological analyser, which, according to the information in the report, would be an FST system fed by 105 suffixes and some 700 verbs. The document does not specify whether these ‘‘verbs’’ are only verb roots or also other type of roots, such as nominals or adjectivals, nor how the parser works in detail. According to some information that has been gathered, the parser would be developed using the Perl programming language and would also be able to generate sentences in Mapudüngun.
@@ 1.1.3 A Lexical Study
In her thesis work (2016), Villena states that her research focuses on examining the authenticity and efficacy of the processes employed in Mapudüngun noun formation. Her objective is to develop criteria for the creation of new nouns in this language. In order to accomplish this objective, she has conducted an analysis of the same dictionaries referenced in §1.1.2.
From the initial statement, it could be inferred that an accurate classification of roots would be of use and performed as a fundamental task. However, the abstract fails to mention that the noun formation under investigation is mainly by derivation processes (compounding is also taken into account), that is to say, by the mechanisms this language commonly uses for the purpose of identifying entities [^14] From our point of view, these are not strictly neologisms, but available (morpho-syntactic) strategies that this language has to deal with the naming of things, new or known. For example, it would not be considered an English neologism: ‘Internet surfer’, even though ‘Internet’ is already a neologism, but ‘Internaut’ would be; it is not registered but possible.
. In any case, the subject of neologicity filters is addressed further down in the work, with the assertion that one of the criteria for determining whether a form is neological is the internal structure of the form (ibid: 11). This suggests that the root component of the neologism should be accurately lexically classified.
However, the prospect of a ‘‘roots classification process’’ appears to have been definitively lost after this note: ‘‘In order to confirm the appropriateness of the lexical categories assigned by Augusta, the most adequate procedure would have been to verify their functioning in discourse. However, the size of the corpus to be analysed made this impossible’’ (ibid: note 82, p. 114). –In the following point §2.1 we explain the procedure followed in our work, which actually implies the process mentioned in Villena’s note–.
Moreover, it appears that accurate lexical classification is not of paramount importance in this research study, particularly in the context of composition, as evidenced by the following citation: ‘‘This finding led to the conclusion that the categories of roots involved in compounding are not crucial in determining the structure of Mapudüngun compounds. Instead, the type of semantic condition maintained between them is deemed to be more significant’’ [^15] As Villena (2016) notes, this is a conclusion of Baker and Fasola (2009) that is supported by the observation of a similarity in the order of the V-N and N-N compounds.
(ibid: 63).
@@
Example E.03.a ((6) b., p.: 53 in the source) shows küdaw- as a verbal root being nominalised by the suffix -fe, which the author says that combines with verbal bases to generate nouns that denote a characteristic agent. But in the dictionaries consulted on the CORLEXIM website (Chandía, 2014) the following entries are found:
  1. küdaw s. The work. (Augusta, F. de (2016a) map-cas)
  2. küdaw N. work. (Smeets, Ineke (2018) map-ing)
  3. küdaw work. (Febrés, A. (1846a) map-cas)
  4. absorto adj. [To be] ngoymaduamkülen (intr.); ngoymakonkülen kiñe küdaw mew. (Augusta, F. de (2016b) cas-map)
@@
As demonstrated above, three authors identify the form as a noun; de Augusta provides an example of the use of the word ‘absorto’ in Mapudüngun (being engrossed in a job), where küdaw is used as a noun without any suffixes, which is in contrast to the classification used by Villena. This example is one of many such instances that can be found in her study.
@@
E.03.b ((92) g., p.: 170 in the source) has been identified as an example of sintagmation, and it is said to represent the Prep-N-V structure. However, a review of dictionaries reveals that:
@@
  1. pu- Vi. to arrive; also puw-. The distribution of pu(w)- is not entirely clear. puw- is most frequent. I found pu-fu-n ‘I arrived, but’ (RR) and iñche pu-n ‘I arrived’ (JM), and I found puw-küle- (RR) only once; müna ayuwüy iñchiu yu puwel ‘he was very glad that we had arrived’; küpa pulen liwen ‘I want to arrive early’; petu puwlaymi küdawmew ‘you are too young to work’; (Smeets, Ineke (2018) map-ing)
@@
Literally translated, E.03.b would mean something like: ‘arrived at the ground as he was going’ or ‘reached the ground on his way’. Perhaps there has been some confusion with another homophonic particle which is a prepositive and which de Augusta and Febrés, among others, define as follows:
@@
  1. pu- artic. pref. pluralises the names of people, animals and some things. pu wentru ‘the men’. pu wangül‘en ‘the stars’. See Augusta (1903, p. 15, 1). (Augusta, F. de (2016a) map-cas)
@@
  2. pu- (part. bef. nouns) denotes plurality; item, at or in; by; under; e.g. pu ovida, sheep (pl.); pu ruka, at home, in the house; pu liwen, in the morning; pu minue, under the ground. (Febrés, A. (1846a) map-cas)
@@
In addition, there is a high probability that the translocative suffix -pu- derives from the same verb, pu-. The affix indicates that the denoted event takes place at a location away from the speaker. This function has also been described by Febrés –and others–:
@@
  1. pu, rpu (interp. to verbs) is to carry out the action of the verb by going from here to there. , v. gr. pepuavin or perpuavin, I’ll drop by to see him; pipuavin or pirpuavin, I’ll pass on to say to him. (Febrés, A. (1846a) map-cas)

### v3

@@ Abstract
After developing a computational system for morphological analysis of the Mapuche language, and evaluating it with texts from various authors and styles, it became necessary to verify the linguistic assumptions of the source used as the basis for implementing this tool. In the present work, the primary focus is on the lexical category classification of Mapudüngun roots recognised as verbal in the source utilised for the development of the morphological analysis system. The results of this lexical category revision directly benefit the computational analyser, as they are implemented as soon as they are verified. Additionally, it is hoped that these results will help clarify some uncertainties about lexical categories in the Mapuche language. This work addresses a preliminary task to identify the valency of true verbal roots, the results of which will be presented in a subsequent work that complements this article.
?abstractname?
@@
Tras haber desarrollado un sistema computacional para el análisis morfológico de la lengua Mapuche y de evaluarlo con textos de diversos autores y estilos, se ha visto la necesidad de verificar los postulados lingüísticos de la fuente utilizada como base para implementar esta herramienta. En el presente trabajo, el enfoque principal es la clasificación de la categoría léxica de las raíces del Mapudüngun reconocidas como verbales en la fuente utilizada para el desarrollo del sistema de análisis morfológico. Los resultados de esta revisión categorial benefician directamente al analizador computacional, ya que se implementan tan pronto como se verifican. Además, se espera que estos resultados ayuden a aclarar algunas incertidumbres sobre las categorías léxicas en la lengua Mapuche. Este trabajo aborda una tarea preliminar a la identificación de la valencia de las raíces verbales, cuyos resultados se presentarán en un trabajo posterior que complementa este artículo.
?abstractname?
Després d’haver desenvolupat un sistema computacional d’anàlisi morfològica de la llengua Mapuche, i d’avaluar-lo amb textos de diversos autors i estils, es va veure la necessitat de verificar els postulats lingüístics de la font utilitzada com a base per implementar aquesta eina. En el present treball, l’enfocament principal és la classificació de la categoria lèxica de les arrels del Mapudüngun reconegudes com a verbals en la font utilitzada per al desenvolupament del sistema d’anàlisi morfològica. Els resultats d’aquesta revisió categorial beneficien directament l’analitzador computacional, ja que s’implementen tan bon punt es verifiquen. A més, s’espera que aquests resultats ajudin a aclarir algunes incerteses sobre les categories lèxiques en la llengua Mapuche. Aquest treball aborda una tasca preliminar a la identificació de la valència de les arrels verbals, els resultats de la qual es presentaran en un treball posterior que complementa aquest article.
@@
Keywords: Mapuche; Mapudüngun; FST; finite state transducer; morphology; morphosyntax; affixes; suffixes; morphemes; verbs; valency; verbal roots; parts of speech
@@
Palabras clave: Mapuche; Mapudüngun; FST; transductor de estados finitos; morfología; morfosintaxis; afijos; sufijos; morfemas; verbos; valencia; raíces verbales; categorías gramaticales; categorías léxicas
@@
Paraules clau: Maputxe; Mapudüngun; FST; transductor d’estats finits; morfologia; morfosintaxi; afixos; sufixos; morfemes; verbs; valència; arrels verbals; categories gramaticals; categories lèxiques
@@ 1   Introduction
This article details the methods used for the reclassification of the lexical roots involved in Mapuche verb themes and presents the results obtained from this process. This task is essential to accurately determine the valency (in a subsequent work, Chandía, 2026) of roots that are genuinely verbal. Without this step, there is a risk of incorrectly attributing verbal properties to roots that do not belong to this lexical category.
@@
The task is described as a reclassification because it involves revisiting the lexical category classification of roots found in Mapudüngun verb themes (see §2.1) which was originally established by the anthropologist Ineke Smeets in her work on the grammatical description of the Mapuche language, published in 2008 with the title “A Grammar of Mapuche”.
@@
This research, like others conducted within the Düngupeyüm [^1] Düngupeyüm is the name of the computational morphological analysis system that has been implemented; the name derives from düngu- ‘language, speech, concern’, -pe suffix indicating physical or temporal proximity, -yü suffix indicating a habitual aspect feature, and the affix -m, a nominaliser indicating place or instrument. The translation would be something like ‘instrument always used for language’, i.e., ‘language tool’. The system is publicly available on https://www.chandia.net/dungupeyum/ project, relies on the grammatical description provided by Smeets (2008). Our goal is to enhance the computational morphological analysis system that we have developed using finite-state transducers (Chandía, 2012, 2021, 2022), which initially implemented the anthropologist’s work. This focus on the elements and dynamics of the language allows us to refine and expand the capabilities of the system.
@@
For a computational system such as the one described earlier, accurately determining both the lexical category of roots and the valency of verbal roots is crucial for proper analysis. This is because, in Mapudüngun, the addition of suffixes to verbal forms, particularly aspectual ones, depends on the verbal valency or its modification during word formation. Consequently, verbal valency plays a key role in analysing the form to identify the implicated suffixes. Therefore, it was necessary to verify the lexical category of roots involved in the formation of verbal themes, as Mapudüngun naturally verbalises roots from other lexical categories, including nouns, adjectives, adverbs, and even some pronouns and numerals. The aim is to accurately recognise the valency of roots that are genuinely verbal, avoiding misclassification of roots belonging to other categories. This document focuses on the task of lexical classification, while the classification of valency will be addressed in a subsequent study (Chandía, 2026).
@@
During the extensive period dedicated to developing the morphological analysis tool, several observations made by Smeets (2008) needed to be verified, and some required modification. She provided a grammatical description of Mapudüngun based on a group of four informants [^2] As asserted by Smeets in email correspondence, the number may be increased by four or five, which where sporadic informants, thus bringing the total to approximately ten.. As a result, her work does not fully capture the entirety of the Mapuche language or all the phenomena present within it.
@@
The main contribution of Smeets (2008)’ “A Grammar of Mapuche”, which is particularly valuable for the development of a computational morphological analysis system, is the identification of a prevalent or fixed position for approximately one hundred suffixes that can appear in Mapuche verbal forms. Additionally, she provides the morphotactics [^3] Morphotactics refers to the rules governing the allowed and prohibited combinations of morphemes (both among themselves and with roots) in a language. These rules influence certain forms within the language. For example, in a Mapuche nominal form, morphemes indicating verbal mode cannot appear unless the nominal form has been verbalised. Similarly, in a verbal form, a nominaliser affix and a grammatical person morpheme cannot co-occur. Morphotactic rules can be influenced by grammatical, semantic and morphological factors., or combinatorial rules, for most of these suffixes.
@@
Smeets (2008) states that suffixes appear in relatively fixed positions along the verbal form. Based on their placement relative to other suffixes and their function, these verbal affixes are organised into thirty-six numbered slots, arranged from right to left. Slot one is at the far right end of the form, while slot thirty-six is next to the root on the left. Certain slots contain suffixes that are mutually exclusive for grammatical or semantic reasons.
@@
E.01 10 suffixes Mapuche verb nünieñmarputueyiñmu
  1. nü      -nie   -ñma   -r    -pu    -tu   -e    -y    -ø   -iñ   -mu (Smeets, 2008: 443; 76)
    tv.take +prps +io +itr +loc +re +inv +ind +1 +pl +3a
    ‘they continued to take it away from us’; Literal: ‘there they went continuously taking it away from us’
Minimal intransitive verb form of 2nd person plural
S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . X X X . . .
Minimal transitive verb form of 2nd → 1st persons, plural
S . . . . . . . . . . . X . . . . . . . . . . . . . . . . . . X ø X . . .
Representation of E.01
S . . . . X . . . . . X . . . . . . . X X X . . . . . . . . . X . X ø X X
?figurename? 1: Graphical representation of the slots in Mapuche verbal forms (root and suffixes)
@@
In the Mapuche language, a verbal theme [^4] To simplify, the term “verbal theme” refers to any root, whether simple or complex, to which suffixes are added to create a complete verbal predication. A simple “verbal theme” consists of a single root, while a complex “verbal theme” may involve two roots, a root with some suffixes, or a combination of these elements. Although it could be argued that these are more accurately described as lemmas, they are all referred to as “verbal theme” for the sake of simplicity (Chandía, 2021: 3). can consist of one, two, or occasionally three roots from different lexical categories, which may combine with each other and sometimes with certain suffixes. A verbal theme can be simple, containing just one root from any lexical category (E.02a, E.02b, E.02c), or complex, with several possible configurations:
  • A root combined with one or more suffixes (E.02d), sometimes up to three suffixes.
@@
    pichi      -ka     -el (Smeets, 2008: 76; 16)
    aj.little +cont +ovn
@@
    tripa     -y    -ø (Smeets, 2008: 67; 51)
    iv.leave +ind +3
@@
    iv.come +ca +ind1sg
    ‘I brought’; Literal: ‘I made come’
@@ 1.1.1   Dictionaries
To date, the most reliable collections of words classified by their lexical type are the dictionaries published throughout the relatively short written history of the Mapudüngun language, particularly the following three [^5] All three dictionaries are accessible on the CORLEXIM site: https://www.chandia.net/corlexim (Chandía, 2014).:
@@
    Havestadt’s dictionary is a fourth lexical source occasionally consulted in cases of doubt. Published in 1777, it provides definitions in Latin [^6] Resource available for consultation via the following link: https://benmolineaux.github.io/bookshelf/1777HAVE-Vocab.html. An alternative Spanish version is also available at the following URL: https://benmolineaux.github.io/bookshelf/1777Havestadt-ParsQuarta-MAPU-SPA.html (Martinez and Carmona, 2023). and can be a valuable resource, particularly when dealing with variations in forms.
@@ 1.1.2   Computational Approach
AVENUE was the name of this initiative, which produced a document entitled “Informe Final” (Álvarez et al., 2005) summarising the project’s achievements and the works still to be done. When it started in 1999, it was known as NICE (Native-Language Interpretation and Communication Environment), later Avenue. Its goals were to create a simultaneous translation system and to contribute to the preservation and development of Indo-American languages [^7] Avenue worked with three indigenous languages, setting up teams in three countries: Aymara in Bolivia, Quechua in Peru and Mapudüngun in Chile.. They were committed to creating their own alphabet for the language, transcribing texts according to this alphabet, extracting grammatical rules from the texts, compiling a parallel corpus of Mapudüngun-Spanish, also a glossary, developing a spell-checker of Mapudüngun for OpenOffice, and building a prototype of an automatic translator of Mapudüngun-Spanish. The report mentioned above states that these tasks were carried out at different levels of completion. The section dealing with the automatic translator mentions that it is an example-based machine translation system [^8] Example-based machine translation, or EBMT, works by searching its database for full translations or, more commonly, parts of sentences that have previously been translated by human translators.. At the time of writing, however, researchers say they are working on a transfer-based machine translation system [^9] In a Transfer-Based Machine Translation (TBMT) system, the original text is first analysed morphologically and syntactically, resulting in a shallow syntactic representation. This representation is then transformed into a more abstract one that emphasises aspects relevant to the translation process and ignores other information. The transfer process converts this latter representation, still linked to the source language, into a representation at the same level of abstraction but linked to the target language. These two representations are called normalised or intermediate representations. From here the process is reversed: the syntactic components generate a representation of the text and finally the translation is generated. in which the basic constructions of Mapudüngun have been recorded: simple sentences with intransitive and transitive verbs, noun phrases with determiners and modifiers, verb phrases with different temporal and aspectual specifications, passive verb phrases, and verb phrases with inverse agreement [^10] The inversion or reverse form in Mapudüngun is one of the possible expressions of the transitive verb (the others are direct and passive constructions) in which there is a third person acting upon a first or second person, this is expressed by the suffix -e before the mood suffix, and the suffix -(m)ew at the end, provided that the first or second person is a patient; i.e. the order of the constituents, which in Mapudüngun are suffixes, is not changed, but the detailed suffixes are added and combined to change the direction of the action of the verb (Zúñiga, 2006: 114).. Also relevant is the section dealing with the morphological analyser, which, according to the information in the report, would be an FST system fed by 105 suffixes and some 700 verbs. The document does not specify whether these “verbs” are only verb roots or also other type of roots, such as nominals or adjectivals, nor how the parser works in detail. According to some information that has been gathered, the parser would be developed using the Perl programming language and would also be able to generate sentences in Mapudüngun.
@@ 1.1.3   A Lexical Study
In her thesis work Villena (2016) states that her research focuses on examining the authenticity and efficacy of the processes employed in Mapudüngun noun formation. Her objective is to develop criteria for the creation of new nouns in this language. In order to accomplish this objective, she has conducted an analysis of the same dictionaries referenced in §1.1.1.
From the initial statement, it could be inferred that an accurate classification of roots would be of use and performed as a fundamental task. However, the abstract fails to mention that the noun formation under investigation is mainly by derivation processes (compounding is also taken into account), that is to say, by the mechanisms this language commonly uses for the purpose of identifying entities [^11] From our point of view, these are not strictly neologisms, but available (morpho-syntactic) strategies that this language has to deal with the naming of things, new or known. For example, it would not be considered an English neologism: ‘Internet surfer’, even though ‘Internet’ is already a neologism, but ‘Internaut’ would be; it is not registered but possible.. In any case, the subject of neologicity filters is addressed further down in the work, with the assertion that one of the criteria for determining whether a form is neological is the internal structure of the form (ibid: 11). This suggests that the root component of the neologism should be accurately lexically classified.
However, the prospect of a “roots classification process” appears to have been definitively lost after this note: “In order to confirm the appropriateness of the lexical categories assigned by Augusta, the most adequate procedure would have been to verify their functioning in discourse. However, the size of the corpus to be analysed made this impossible” (ibid: note 82, p. 114). –In the following point §2.1 we explain the procedure followed in our work, which actually implies the process mentioned in Villena (2016)’s note–.
Moreover, it appears that accurate lexical classification is not of paramount importance in this research study, particularly in the context of composition, as evidenced by the following citation: “This finding led to the conclusion that the categories of roots involved in compounding are not crucial in determining the structure of Mapudüngun compounds. Instead, the type of semantic condition maintained between them is deemed to be more significant” [^12] As Villena (2016) notes, this is a conclusion of Baker and Fasola (2009) that is supported by the observation of a similarity in the order of the V-N and N-N compounds. (ibid: 63).
@@
Example E.03a ((6) b., p.: 53 in the source) shows küdaw- as a verbal root being nominalised by the suffix -fe, which the author says that combines with verbal bases to generate nouns that denote a characteristic agent. But in the dictionaries consulted on the CORLEXIM website (Chandía, 2014) the following entries are found:
  1. 1.
    küdaw s. The work. (de Augusta map-cas in Chandía, 2014)
  2. 2.
    küdaw N. work. (Smeets map-ing in Chandía, 2014)
  3. 3.
    küdaw work. (Febrés map-cas in Chandía, 2014)
  4. 4.
    (to be) engrossed adj. ngoymaduamkülen; (de Augusta cas-map in Chandía, 2014)
    ngoymakonkülen kiñe küdaw mew ‘to be engrossed in a task".
@@
As demonstrated above, three authors identify the form as a noun; de Augusta provides an example of the use of the word ‘absorto’ in Mapudüngun (being engrossed in a job), where küdaw is used as a noun without any suffixes, which is in contrast to the classification used by Villena (2016). This example is one of many such instances that can be found in her study.
@@
E.03b ((92) g., p.: 170 in the source) has been identified as an example of sintagmation, and it is said to represent the Prep-N-V structure. However, a review of dictionaries reveals that:
@@
pu- Vi. to arrive; also puw-. The distribution of pu(w)- is not entirely clear. puw- is most frequent. I found pu-fu-n ‘I arrived, but’ (RR) and iñche pu-n ‘I arrived’ (JM), and I found puw-küle- (RR) only once; müna ayuwüy iñchiu yu puwel ‘he was very glad that we had arrived’; küpa pulen liwen ‘I want to arrive early’; petu puwlaymi küdawmew ‘you are too young to work’ (Smeets map-ing in Chandía, 2014.)
@@
Literally translated, E.03b would mean something like: ‘arrived at the ground as he was going’ or ‘reached the ground on his way’. Perhaps there has been some confusion with another homophonic particle which is a prepositive and which de Augusta and Febrés, among others, define as follows:
@@
pu artic. pref. pluralises the names of people, animals and some things. pu wentru ‘the men’. pu wangülen ‘the stars’. See Augusta (1903, p. 15, 1) (de Augusta map-cas in Chandía, 2014.)
@@
pu (part. bef. nouns) denotes plurality; item, at or in; by; under; e.g. pu ofida, sheep (pl.); pu ruka, at home, in the house; pu liwen, in the morning; pu minue, under the ground (Febrés map-cas in Chandía, 2014)
@@
In addition, there is a high probability that the translocative suffix -pu derives from the same verb, -pu. The affix indicates that the denoted event takes place at a location away from the speaker. This function has also been described by Febrés –and others–:
@@
pu, rpu (interp. to verbs) is to carry out the action of the verb by going from here to there. , v. gr. pepuafin or perpuafin, I’ll drop by to see him; pipuafin or pirpuafin, I’ll pass on to say to him (Febrés map-cas in Chandía, 2014)